Hi,
I need help with a problem. I can figure it out, but need to set it up in an algebra equation. Here it is: Irving School, Hoover School, and West School participated in a clean-up campaign to collect scrap aluminum. Irving School collected 40% more scrap aluminum than Hoover School. Hoover School collected 25% more than West School. If the total collected by the 3 schools was 560 kg, how much did Irving School collect?

Answer by checkley71(8403)

I=1.4X1.25X west
H=1.25 west
W=west
therefore 1.4x1.25W + 1.25W + W = 560
1.75W +1.25W + W = 560
3W=560 or W=140 I=1.75X140 OR I=175 AND H=1.25X140 OR H=245
PROOF 140+175+245=560